help please
solve the rational inequality (x-2)/(x+3) is less than or equal to 0
is it: (-3,2] (-oo,-3) (-oo,-3)U{2,oo) or [2,oo)
i am very confused

Answer by jai_kos(139)

Given (x-2)/(x+3) is less than or equal to 0
(x-2) /(x-3) < = 0
multiply the above equation by (x -3), we get
(x-2) < = 0
x -2 < =0
Add 2 on both sides, we get
x < = 2
Therefore this in the interval form is given by
(-infinty ,2]